OpenStudy (anonymous):

A car is traveling at 20.0 m/s, and the driver sees a traffic light turn red. After 0.530 s (the reaction time), the driver applies the brakes, and the car decelerates at 7.00 m/s2. What is the stopping distance of the car, as measured from the point where the driver first sees the red light?

OpenStudy (anonymous):

d=vt+1/2at^2 t= 0.530 s v=20.0 m/s a=7.00 m/s2 solve for d

OpenStudy (anonymous):

Yes I tried this but the answer in the back of the book is 39.2m I get 9.62m.

OpenStudy (anonymous):

his reaction time is .53 seconds, meaning for .53 seconds, he's still travelling at 20 m/s now after .53 seconds, he starts to decelerate 20-7t=0 solve for t, this is the amount of time in order to reach 0 velocity or in other words, to stop then plug that into the distance x=1/2 (-7)t^2+20t+x0 x0 is the same as the distance travelled during the reaction time (.53*20)

hartnn (hartnn):

for 0.53 sec car travels with 20m/s--->distance covered = 10.6 m now since it decelerates, a=-7 u=initial vel = 20, v =0(car stops) v^2=u^2+2as 0=400-2*7*s s=28.57 total distance = 10.6+28.57 = 39.17m

OpenStudy (anonymous):

first calculate the distance car travels during response time D1 = speed * time after that it de-accelerate by 7m/s^2 v^2 - u^2 = 2a*D2 put the values of v,u,a net distance is D1 + D2

OpenStudy (anonymous):

thank you ALL for your help

hartnn (hartnn):

welcome :) good question.